### v1

# Einstein’s basement - a model for dark matter and an expanding universe?

Abstract
We present a model treating the energy-momentum relation of relativistic particles as the upper branch of a generalized energy-momentum relation emerging from a forbidden crossing between the constant energy of a massive particle and the photon line. The lower branch, a regime dubbed as Einstein’s basement, gives rise to particles with different kinematics that is analysed in the low-velocity limit. Allowing for gravitational interaction between those particles, and between particles of both branches, we discuss whether particles in Einstein’s basement might be suitable dark matter candidates. Moreover, the special dynamics of mixed systems leads to an expansion mechanism for regular matter in space. Tests of the presented model by astronomical observations are suggested.
Introduction
The large scale dynamics of the Universe by regular matter is understood only if we allow for hypothetical dark matter and dark energy [1]. Dark matter [2] is a well established concept, allowing to describe a huge number of observations in cosmology [3, 4], besides other explanations such as “Modified Newtonian Dynamics (MOND)” [5, 6]. Whereas regular matter is successfully described by the standard model of particles [7], the origin of dark matter is not yet clear. A number of theoretical hypotheses has been offered for the origin of dark matter like primordial black holes [8], weakly interacting massive particles (WIMPs) [9], axion-like particles [10], massive compact halo objects (MACHOs) [11], dark photons [12], erebons [13], Kaluza-Klein particles [14] or others. Another attempt to explain dark matter and dark energy was based on the assumption of negative masses [15]. More than fifty direct or indirect experimental approaches for detection of the WIMPs and axions are under way [16] but up to now none of the proposed candidate particles have been found to explain dark matter to the extent to which it is present in the universe [17, 18].
Any attempt to solve the presently unexplained mysteries is stepping back from the known physics in one point or another. A systematic proof of the assumptions of standard model physics therefore is essential for constraining New Physics, and the degree of how strong known principles of physics can be violated. Here, we explore a possible extension where we will relax our restrictions on what is considered to be a particle.
We rely on a more general view of a “particle” as any entity with a well defined energy momentum dispersion relation (EMR). Among those are also so-called “quasi particles” or “collective excitations” where a particular interaction dresses the bare particles with modified properties. This fruitful concept is used, e.g., in Fermi liquids [19], phonons or polaritons [20, 21, 22] in solid state physics, but may also apply to the Standard Model where particles, normally considered as fundamental, are treated as excitations of their underlying quantum fields.
The paper is organized as follows. First, we find besides the well known relativistic EMR a second one whose energy is below the former, marking a region that is dubbed henceforth as “Einstein’s basement”. Then, as a logical first step, we apply classical mechanics to the new EMR and derive the corresponding equations of motion analogue to the mechanics of special relativity and Newtonian physics. We then explore whether such quasi particles might represent candidates for dark matter and give examples of possibly observable phenomena.
In special relativity a free particle has the well-known EMR
$E_{SR}(p)=\sqrt{(mc^{2})^{2}+{\mathbf{p}}^{2}c^{2}}$ (1)
with $m$ the rest mass of the particle, $c$ the speed of light in vacuum and $p=|{\mathbf{p}}|$ the modulus of the momentum of the particle [23, 24]. This EMR is (the only one) intimately associated with the geometry of Minkowski spacetime governed by Lorentz symmetry. Its properties have been tested to its best over the last century. This EMR is shown in Fig. 1 as the upper two curves for two different masses $m_{1}$ and $m_{2}$. For low velocities each curve starts with the respective rest mass energy $m_{i}c^{2}$ before it turns into a linear behavior for large kinetic energies.
I The model
Our model starts with the notion that the EMR of special relativity represents the upper branch of an avoided crossing between the EMR of a constant matter bare state $E_{b1}=mc^{2}$ and the EMR of a photon bare state $E_{b2}=pc$. In a simple form an avoided crossing can be described by a matrix $\cal{A}$ that couples these two bare states with the coupling $F(p)$ as
[…]
Diagonalizing this matrix leads to the dressed state $E(p)$ with the characteristic equation
$E(p)^{2}-(pc+mc^{2})E(p)+pcmc^{2}-F(p,m)^{2}=0\,.$ (3)
Assuming the first solution to be $E_{SR}(p)=\sqrt{p^{2}c^{2}+m^{2}c^{4}}$ the second solution is
$E_{EB}(p)$ $=$ $mc^{2}+pc-\sqrt{p^{2}c^{2}+m^{2}c^{4}}\,,$ (4)
[…]
Figure 1: The relativistic energy momentum relations of two free particles (upper two curves) with rest masses $m_{1},m_{2}$ as function of the particles’ momenta. They approach the respective rest energies $E=m_{i}c^{2}$ (horizontal lines) for low momenta and the energy-momentum relation of a photon (dash-dotted line) for high momenta. Furthermore, for each $m_{i}$ a second branch is shown, which, in contrast to the preceding case, asymptotically approaches the photonic line at low momenta and the respective mass lines at high momenta. The gray area represents what we refer to as Einstein’s basement in this paper.
The new EMR of Eq. (4) is displayed also in Fig. 1 for the same two masses $m_{1},m_{2}$. For small momenta, they approach the photon line whereas they cling to their rest masses for large momenta.
[…]
Having assumed, that energy is a valid concept in Einstein’s basement - in the sense of a first integral of motion - we will also assume that Hamiltonian mechanics [28] with the Hamiltonian ${\cal{H}}={\cal{H}}({\mathbf{q}},{\mathbf{p}},t)$, and the canonical coordinates, position ${\mathbf{q}}$ and momentum ${\mathbf{p}}$, is suitable to describe the physics in both branches.
The velocity $\mathbf{v}$ of a particle is given by the Hamilton equation
$\mathbf{v}=\dot{\mathbf{q}}=\nabla_{\mathbf{p}}\cal{H}\,.$ (5)
Hence, for ${\cal{H}}=E_{SR}$ we obtain
$\mathbf{v}_{SR}=\frac{pc^{2}}{\sqrt{(mc^{2})^{2}+{p}^{2}c^{2}}}\hat{\mathbf{v}}\,,$ (6)
where $\hat{\mathbf{v}}$ is the velocity unit vector. In contrast, for the lower branch ${\cal{H}}=E_{EB}$ we get
$\mathbf{v}_{EB}$ $=$ $\left(c-\frac{pc^{2}}{\sqrt{(mc^{2})^{2}+p^{2}c^{2}}}\right)\hat{\mathbf{v}}\,.$ (7)
Figure 2: Velocity of the particle as a function of the momentum $p$ according to Eq. (6) and (7) for $E_{SR}$ (full line) and $E_{EB}$ (dashed line), respectively.
Thus, we find that the velocity of a particle behaves differently in both cases. In special relativity $|\mathbf{v}|=v$ depends linearly on the momentum for $v\ll c$ and approaches the velocity of light for asymptotically high momenta. In Einstein’s basement $v$ decreases from $c$ for low momenta and goes to zero for $p\to\infty$. Both behaviors are illustrated in Fig. 2.
From Eq. (6) one derives the well known momentum of a relativistic particle
$\mathbf{p}_{SR}=\frac{mv}{\sqrt{1-\left(\frac{v}{c}\right)^{2}}}\hat{\mathbf{v}}.$ (8)
Doing the same for the lower branch, we obtain
$\mathbf{p}_{EB}=\frac{m(c-v)}{\sqrt{1-(\frac{c-v}{c})^{2}}}\hat{\mathbf{v}}.$ (9)
We find that both momenta have a similar form, while in Eq. (9) the velocity $v$ is replaced by $c-v$ in comparison to the relativistic case. Thus, we observe that both branches respect $c$ as an ultimate limit for the motion of particles, while the roles of $v=0$ and $v=c$ are interchanged. In the relativistic case, the special relation of the momentum (8) and the energy (1), gives rise to the class of Lorentz transformations, keeping $\sqrt{E_{SR}^{2}-c^{2}\mathbf{p}_{SR}^{2}}=mc^{2}$ invariant. However, this does not hold for the lower branch, where no such transformation can be found. In particular, $\sqrt{E_{EB}^{2}-c^{2}\mathbf{p}_{EB}^{2}}$ is not Lorentz invariant by itself, i.e., has a different value in every Lorentz frame, giving up this principle of relativity in the lower branch solely. We will see that this feature will also translate into the equations of motion, when we investigate the dynamics of the newly introduced particles.
To study the kinematics of particles in Einstein’s basement, we want to construct the counterpart of the well known relativistic Lagrangian
$\mathcal{L}_{SR}$ $=$ $-mc^{2}\sqrt{1-\left(\frac{v}{c}\right)^{2}}$
$=$ $-mc^{2}+\frac{m}{2}\,v^{2}-\dots\,.$
for a free particle, where an expansion for small velocities $v\ll c$ reveals the well-known Newtonian kinetic energy term. Keeping in mind that the relation
$\mathbf{p}_{EB}=\frac{\partial{\cal L}_{EB}}{\partial\mathbf{v}}$ (11)
holds, we integrate Eq. (9) to obtain the Lagrangian
${\cal L}_{EB}$ $=$ $-{m}c^{2}+{m}c^{2}\sqrt{1-\left(\frac{c-v}{c}\right)^{2}}$
$=$ $-{m}c^{2}+\frac{{m}}{2}(2c)^{3/2}\sqrt{v}-\dots$
for a free particle in Einstein’s basement, where the integration constant was chosen as $-{m}c^{2}$ to obtain the Lagrangian for a particle at rest. In analogy to the second line of Eq. (I), we obtain a kinetic energy term in the low-velocity limit. Remarkably, also in this limit, the speed of light $c$ remains as a parameter in the system, affecting the motion of such particles. In the next section, we will analyze this expression and discuss the resulting dynamics of particles, which qualitatively differs from the Newtonian case in many aspects.
II New dynamics in Einstein’s basement
In the following, we have a closer look at the dynamics in Einstein’s basement by using the free particle Lagrangian Eq. (I) in the low-velocity limit and complement it by an additional potential depending on the particle’s position in space
~$\tilde{\mathcal{L}}=\mathcal{L}_{EB}-V(\mathbf{q})\,.$ (13)
The corresponding Euler-Lagrange equations
$\frac{d}{dt}\frac{\partial{\cal L}_{EB}}{\partial\mathbf{v}}=-\frac{\partial V(\mathbf{q})}{\partial{\mathbf{q}}}$ (14)
lead to the equations of motion in the form
$m\,\mathbf{a}-\frac{3m(\mathbf{a}\cdot\mathbf{v})}{2v^{2}}\,\mathbf{v}=\frac{1}{2}\left(\frac{2v}{c}\right)^{3/2}\mathbf{F}\,,$ (15)
where $\mathbf{a}=\ddot{\mathbf{q}}$ is the particle’s acceleration and $\mathbf{F}$ the force resulting from the potential. We find that the time derivative of the momentum $\frac{d}{dt}\mathbf{p}_{EB}\not=m\,\mathbf{a}$, does not coincide with the acceleration, as it is the case in Newtonian physics. In contrast, $m\,\mathbf{a}$ is complemented by an additional term, which depends on the particle’s velocity $\mathbf{v}$. By rearranging Eq. (15) in the form
${m}\,{\mathbf{a}}=\frac{1}{2}\left(\frac{2v}{c}\right)^{3/2}\left(\mathbf{F}-3\frac{(\mathbf{F}\cdot\mathbf{v})}{v^{2}}\,\mathbf{v}\right)$ (16)
the equations of motion can be directly compared to Newton’s second law. For $\mathbf{F}=0$ we obtain the result $\mathbf{a}=0$ as in the Newtonian case, leading to a uniform motion of the particle. In the presence of external forces, however, the dynamics of the particle differs from Newton’s second law in several aspects. We find that the effect of the external force is drastically reduced for slow particles due to the overall-prefactor, depending on $|\mathbf{v}|=v\ll c$. However, the orientation of $\mathbf{v}$ with respect to $\mathbf{F}$ affects the term in the second brackets. Due to that, the particle’s response to the force can be positive or negative, depending on the velocity vector. For instance, in the special case when the force and the velocity are parallel (or anti-parallel), we obtain an overall negative sign
${m}\,{\mathbf{a}}=-\left(\frac{2v}{c}\right)^{3/2}\mathbf{F}\,,$ (17)
such that in this case commonly attractive forces, act repulsive on a “basement particle”. We stress, that this repulsive force has its origin in the particle dynamics and must not be confused with the assumption of a negative particle mass, as it was assumed in dark energy considerations by A. Einstein [29], or more recently by J. S. Farnes [15].
Having disclosed the unconventional behavior of particles in the lower branch of Fig. 1 with respect to external forces, it is straightforward to investigate the interaction of basement particles with each other and with common matter. In the following, we will consider a purely gravitational interaction between two classical particles, two particles in Einstein’s basement, and between one of each kind.
II.1 Dynamics of two classical particles
In order to study the gravitational interaction of the newly introduced particles, we first recall the central features of the Keplerian motion of two regular particles of masses $m_{1}$ and $m_{2}$. The corresponding Lagrangian is given by
$\mathcal{L}_{SR-SR}=\frac{m_{1}}{2}v_{1}^{2}+\frac{m_{2}}{2}v_{2}^{2}+\frac{Gm_{1}m_{2}}{|\mathbf{q}_{1}-\mathbf{q}_{2}|}$ (18)
By utilizing the Euler-Lagrange equations and the definition of the relative and center of mass coordinates
$\mathbf{r}$ $=$ $\mathbf{q}_{1}-\mathbf{q}_{2}$ (19)
$\mathbf{s}$ $=$ $\frac{m_{2}}{m_{1}+m_{2}}{\mathbf{q}}_{2}+\frac{m_{1}}{m_{1}+m_{2}}{\mathbf{q}}_{1}\,,$ (20)
the dynamics of the system is described by
$\ddot{\mathbf{r}}$ $=$ $-\frac{G(m_{1}+m_{2})}{|\mathbf{r}|^{3}}\,\,\mathbf{r}$ (21)
$\ddot{\mathbf{s}}$ $=$ $0\,.$ (22)
Therefore, one finds that the relative motion of the particles and their center of mass motion can be described independently. The equation (21) leads to the well-known Kepler orbits while the particle’s center of mass moves with a constant velocity. We will see that neither the first nor the second condition will hold when particles from Einstein’s basement are involved.
II.2 Dynamics of two slow basement particles
In analogy to the latter case we again define the Lagrangian
$\mathcal{L}_{EB-EB}$ $=$ $\frac{m_{1}}{2}(2c)^{3/2}\sqrt{v_{1}}+\frac{m_{2}}{2}(2c)^{3/2}\sqrt{v_{2}}$ (23)
$+$ $\frac{Gm_{1}m_{2}}{|\mathbf{r}|}$
of two gravitationally interacting particles using Eq. (I) in the low-velocity limit. The equation of relative motion for the two particles reads
$\ddot{\mathbf{r}}$ $=$ $-\frac{\sqrt{2}G\left(m_{1}(v_{2}/c)^{3/2}+m_{2}(v_{1}/c)^{3/2}\right)}{|\mathbf{r}|^{3}}\,\,\mathbf{r}$ (24)
$+\,\frac{3\sqrt{2}G}{c^{3/2}|\mathbf{r}|^{3}}\left(\frac{m_{2}(\mathbf{v}_{1}\cdot\mathbf{r})}{\sqrt{v_{1}}}\,\,\mathbf{v}_{1}+\frac{m_{1}(\mathbf{v}_{2}\cdot\mathbf{r})}{\sqrt{v_{2}}}\,\,\mathbf{v}_{2}\right)\!.\,\,\,$
This equation contains a number of important and counter-intuitive features. First, it shows that the motion of the particles does not only depend on their relative position $\mathbf{r}$, but also on their velocities $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ with respect to an absolute space. Second, the first term of Eq. (24) is attractive like in the Newtonian case from Eq. (21), but with a velocity-dependent mass term $\sqrt{2}[m_{1}(v_{1}/c)^{3/2}+m_{2}(v_{2}/c)^{3/2}]$ instead of the total mass $m_{1}+m_{2}$ of the system. Thus, for small velocities $v_{i}\ll c$ the mass effectively entering the gravitational interaction is drastically reduced in comparison to the classical two-body problem. Third, we find the second, also velocity-dependent term in Eq. (24) which is absent in the Newtonian case. It is always repulsive and can be up to three times larger than the attractive term. Consequently, the interplay of these two contributions decides whether the net acceleration between the two particles is positive or negative.
To study the dynamics of two gravitationally interacting basement particles in more detail we consider the three ultimate cases in Fig. 3.
Figure 3: Orbit of a basement particle with mass $m_{2}$ and velocity $\mathbf{v}_{2}$, orbiting another basement particle of mass $m_{1}$ at rest. According to Eq. (24) the moving particle describes a circular orbit in the particular case of $v_{2}=R_{0}\Omega$. Below this value, for $v_{2}<R_{0}\Omega$, the second particle comes to a final stop (at the red cross) within the circular orbit. For $v_{2}>R_{0}\Omega$ the second particle increases its distance and eventually escapes to infinity.
For that purpose we assume the first particle to be at rest with respect to absolute space, i.e., $\mathbf{v}_{1}=0$. Further assuming $\mathbf{v}_{2}\cdot\mathbf{r}=0$, we find that there exists a circular orbit of the second particle with $|\mathbf{r}|=R_{0}=\mathrm{const}$ and a constant frequency $\Omega=\frac{2G^{2}m_{1}^{2}}{c^{3}R_{0}^{3}}$. The motion of the second particle, regardless of its own mass $m_{2}$, only depends on the mass $m_{1}$ of the fixed particle and the distance $R_{0}$. This follows naturally recalling that the momentum of a basement particle with zero velocity is infinite, cf. Eq. (7) and Fig. 2. The circular orbits with velocity $v_{2}=R_{0}\Omega$ are unstable and mark the transition between two regimes of distinct behaviour. If the velocity $v_{2}>R_{0}\Omega$ of the second particle is increased by a very small amount, the particle escapes to infinity, as shown in Fig. 3b. If the velocity $v_{2}<R_{0}\Omega$ of the second particle is reduced, the particle turns into the region of the circular area and ultimately comes to rest at a finite distance, as shown in Fig. 3c. Such a stopping point is also encountered when two basement particles of arbitrary mass move towards each other on a straight line in absolute space.
II.3 Dynamics of a classical and a (slow) basement particle
Here, we want to discuss the behavior of the relative motion and the non-conservation of the center of mass velocity for the gravitational interaction of classical and a basement particle. Beginning with the Lagrangian
$\mathcal{L}_{SR-EB}=\frac{m_{1}}{2}v_{1}^{2}+\frac{m_{2}}{2}(2c)^{3/2}\sqrt{v_{2}}+\frac{Gm_{1}m_{2}}{|\mathbf{r}|}$
and the corresponding Euler-Lagrange equations we find
$\ddot{\mathbf{q}}_{1}$ $=$ $-\frac{Gm_{2}}{|\mathbf{r}|^{3}}\mathbf{r}$ (25)
$\ddot{\mathbf{q}}_{2}$ $=$ $\frac{\sqrt{2}Gm_{1}\left(v_{2}^{2}\,\mathbf{r}-3(\mathbf{v}_{2}\cdot\mathbf{r})\,\mathbf{v}_{2}\right)}{c^{3/2}|\mathbf{r}|^{3}\sqrt{v_{2}}}\,.$ (26)
Other than in the Lagrangian, at the level of the equations of motion we can now easily perform the limit $v_{2}/c\to 0$ and find
$\ddot{\mathbf{r}}=-\frac{Gm_{2}}{|\mathbf{r}|^{3}}\mathbf{r}\qquad$ , $\qquad\ddot{\mathbf{q}}_{2}=0\,,$ (27)
where $\ddot{\mathbf{r}}=\ddot{\mathbf{q}}_{1}$ holds in this case. We find, that the classical particle behaves like a test particle attracted by the gravitational potential of a fixed particle of mass $m_{2}$. Within this approximation the basement particle is still allowed to move with a constant but small velocity $v_{2}\ll c$, such that the total system acts like a classical system with $m_{1}\ll m_{2}$ under Galilean transformation. Other than that, however, the relation between the masses plays no role if the particle of mass $m_{2}$ is a slow basement particle. We see from Fig. 4 that the motion of the particle in Einstein’s basement is virtually not affected by the classical particle whereas the latter one’s motion encircles the first one. Hence, the center of mass velocity of the combined system is not a constant of motion.
Figure 4: Simulated trajectories of a regular particle with mass $m_{2}$ (black) and a slow basement particle with mass $m_{1}$ (blue), interacting gravitationally. The basement particle is almost unaffected by its companion, whereas the regular particle follows the former one in a spiraling motion, see (a) and (b). The relative motion of $m_{2}$ with respect to $m_{1}$, given in (c), shows two periodicities: an elliptic orbit comparable to Kepler motion, and an oscillation of the apsidal position.
III Are particles in Einstein’s basement dark matter candidates?
We are now in the position to investigate whether our model can reproduce the central features of dark matter.
The mass of dark matter.
In our approach we treat the EMR of regular matter as one of the two branches of a generalized energy-momentum relation that emerges from a forbidden crossing of the constant energy of a massive particle and the photon line. This leads to a new zoo of particles, having the same masses as their relativistic counterparts. The emerging new sector could complement the known constituents of the current standard model.
The existence of dark matter was first deduced by Fritz Zwicky [30] from observed discrepancies in the dynamics of galaxy clusters from the laws of celestial mechanics, such as the virial theorem. In contrast to the usual approach to attribute these observations solely to an invisible distribution of matter with regular kinematics, in our model the emerging particles have a different kinematics as in the regular relativistic or Newtonian case. This mechanism can be an alternative to the usual dark matter model.
Photons, and why the basement may be dark.
There is a striking difference between the regular EMR $E_{SR}(p)$ and $E_{EB}(p)$ in Einstein’s basement with respect to the possibility of mass-less particles by setting $mc^{2}=0$. In the first case we recover $E_{SR}(p)=pc$, i.e., the linear EMR of photons. In contrast, for a mass-less particle in Einstein’s basement we find $E_{EB}(p)=0$, such that it never has an energy other than zero. Consequently, there are no photon-like particles in Einstein’s basement. This, a priori, would not prevent that a particle in Einstein’s basement could interact with a regular photon. The interaction between a photon and a particle in Einstein’s basement would require an additional coupling between both branches. Let us look on the probability of such a hypothetical interaction from a different viewpoint. Diabatic electromagnetic transitions are well known to occur in the case of an avoided crossing. Such transitions were modeled long ago by Landau [31], Zener [32], Stückelberg [33] and Majorana [34] and have been tested experimentally in quantum optics [35]. Assuming that diabatic transitions under the influence of one or more of the well known interactions are allowed, they would lead to an increased population in the $E_{EB}$ state which has a lower energy than $E_{SR}$. Consequently, direct transitions between the two regimes, and, hence, the direct interaction between both states must be extremely low. As a result, we conclude that there is not much room for the interaction of photons with particles in Einstein’s basement.
Low coupling and the predominance of dark matter over regular matter
While we can assume that a spontaneous decay of classical particles into the basement is highly suppressed at the current age of the universe, the interaction between both branches may have changed during its different phases. However, the basement as the energetically more favorable state can be an explanation of the current dominance of dark matter in comparison to common matter in the present universe.
Clusters of dark matter.
For regular matter, the formation of large bodies by accumulation of microscopic particles via gravitational attraction is well understood. A possible aggregation of dark matter consisting of particles from Einstein’s basement would be more intricate due to the repulsive and attractive forces in the game.
The velocities of cold dark matter in galaxies and spiral nebula have been determined to be in the range of $200-550$ km/s [36, 37]. As we discussed in Sec. II.3, regular matter is strongly attracted by such slow basement particles, see Eq. (27), while the potentially repulsive forces between the basement particles are suppressed by $(v/c)^{3/2}$, see Eq. (24). Moreover, given the attractive forces between regular matter components, such a mixture could be stable even though the fast constituents of dark matter would boil off.
The preceding arguments lead us to assume that the quasi particles derived from our model of Einstein’s basement can be regarded as possible dark matter candidates.
Having discussed the relation of basement particles and the observed properties of dark matter, we want to emphasize yet another feature of a mixed system of basement particles and regular matter. Under the assumption of classical gravitational interaction, particles in Einstein’s basement - apart from special cases - are repelled by each other, while regular particles follow them regardless of their own mass. This mechanism describes an expansion of a mixed system in space. Whether this can be a source of the observed accelerated expansion of the universe, as it is suggested by the study of type Ia supernovae candles [38, 39], has to be investigated. The modeling of statistically significant ensembles is indispensable to answer this question. It is also necessary to check if our model can explain other observations already described reasonably well by the models of $\Lambda$CDM and MOND at large and small scales, respectively.
IV Summary, Outlook and Conclusions
[…]
Our findings might possibly be tested quantitatively at different scales, for instance by astrophysical observations of stars, nebulae and galaxies and by gravitational lensing effects of regular and dark matter components. A promising candidate for such an investigation could be the so-called bullet cluster [40]. This cluster comprises of two components each with a spearhead dark matter contribution followed by a distribution of hot regular gaseous material. In such a geometry our model asks - in contrast to regular kinematics - for additional acceleration of the two dark matter components away from each other, which might be observable in successive measurements.
It must be checked, whether our model delivers an competitive explanation for the existing measurement data and provides an alternative to the current maps of the dark matter distribution in the galaxies [41, 42]. Revisiting the huge number of existing data and the new data expected from the EUCLID space mission [43] could give a hint whether the presented ideas should be pursued further.
If our model can survive the comparison with astrophysical observations there would be many more options to explore new physics. In this paper we only considered classical particle dynamics and compared our results to those of classical Newtonian physics. However, with the full Hamiltonian (I) at hand, an extension of our analysis towards a comparison with special and general relativity is possible. This is of particular interest, since the concept of relativity does not hold for the sector of Einstein’s basement, while it is still preserved for regular matter in our model. However, considering interactions between particles of both sectors, violations of the relativity principle, such as a breaking of local Lorentz invariance, can be expected also for regular matter.
A field theoretical approach of our model could elucidate the nature of the field whose excitations are the quasi particles, we identified as regular particles and particles from Einstein’s basement. If we stress the similarities of our model with the description of polaritons in a crystal [20, 21, 22], the underlying field would correspond to the crystal lattice in that case. In particular, the relation of the mass or energy density of this hypothetical background field to concepts like the cosmological constant and dark energy has to be investigated.
A consequent extension of the work presented here is to study the given EMRs in a quantum mechanical framework and, ultimately, within the formalism of quantum field theory. Particularly interesting is the regime of small energies and momenta, where all EMRs in Einstein’s basement closely approach each other and the photon line.
Acknowledgements.
We thank Uwe Sterr, Piet Schmidt, Andrey Surzhykov, Eberhard Tiemann and Jun Ye for fruitful discussions and Christopher Mertens for numerical simulations. We acknowledge the support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy—EXC 2123 QuantumFrontiers—390837967. FR thanks QUEST Leibniz Research School for longstanding support.